OpenStudy (anonymous):

WILL GIVE MEDAL Solve the problem. The formula d = rt gives the distance traveled in time t at rate r. If a bicyclist rides for 2.3 hours and travels 41 miles, what was her average speed? A. 9.43 miles per hour B. 5.6 miles per hour C. 17.83 miles per hour D. 38.7 miles per hour

OpenStudy (anonymous):

the answer is obvious I mean you cant go 38.7 mph on a bike or 17.83 mph

OpenStudy (anonymous):

o

OpenStudy (***[isuru]***):

ok... the equation to find the distance is distance = time x rate... which is d = rt if we want to find the rate .. we can write it as rate = distance / time which is r = d/t where d = distance = 41 miles t = time = 2.3 hours so... r= 41/2.3 = 17.83 miles per hour :)